8. Suppose you buy a $100 government bond that is due next year. How much nominal interest will
you receive if inflation is 4 percent over the year and the bond promises a real return of 3 percent?
1
Expert's answer
2021-08-09T15:42:34-0400

Nominal rate of return= 7.12%

Explanation:

Inflation is the increase in the price level. It erodes the value of money. 

Nominal interest is that quoted for investment or loan transactions. It has not been been adjusted for inflation. 

Real interest rate is the amount of interest in terms of the the quantity of good and services that can be purchased. It is the nominal interest rate adjusted for inflation. 

The relationship between inflation, real interest and nominal interest rate is given using the Fishers Effect; 

N = ( (1+R) × (1+F)) - 1

N- nominal rate, R-real rate, F- inflation 

Nominal rate of return =(1.03)× (1.04) - 1 =0.0712

Nominal rate of return = 0.0712 × 100 = 7.12%

Nominal rate of return= 7.12%